OpenStudy (anonymous):

TRUE OR FALSE. A sock is drawn from a drawer containing 9 white socks, 8 blue socks, and 10 brown socks. The probability of getting a brown sock is 42%.

OpenStudy (anonymous):

false

OpenStudy (anonymous):

explanation: total socks=8+9+10=27; number of brown socks=10 so the probability will be 10/27 = \[\approx37%\]

OpenStudy (anonymous):

per cent